STUDY OF DISTORTED SPHERICAL TOPS USING SYMMETRY-ADAPTED TENSORIAL FORMALISM: APPLICATION TO SF_{5}Cl AND SO_{2}F_{2}

\underline{M. ROTGER}, V. BOUDON, M. LOETE, Laboratoire de Physique de l’Université de Bourgogne – CNRS UMR 5027, 9 Av. A. Savary, BP 47 870, F-21078 DIJON Cedex, FRANCE.

The techniques of symmetry-adapted tensorial formalism and of vibrational extrapolation developed since many years have proved their efficiency for the spectroscopy of spherical-top molecules (CH_{4}, SF_{6}, …)^{a}. We have recently extended these methods to the case of “distorted spherical-tops”, i.e. XY_{4} or XY_{6} type molecules for which one or two of the Y ligands has been substituted by another atom. To do this, we start from the O(3) ⊃ T_{d} or O(3) ⊃ O_{h} formalism and perform a reorientation into the subgroup C_{2v} or C_{4v}. We have first considered the case of XY_{5}Z (C_{4v}) molecules^{b}^{,}^{c}. This model has been used to perform the analysis of the ν_{1}/ν_{8} dyad of SF_{5}^{35}Cl^{d} A set of programs for spectrum calculations and fits using these methods has been developed^{e} and is made freely available through the World Wide Web^{f}. We also plan to use these programs for the spectroscopy of SF_{5}Br or IOF_{5}, for instance. We have also undertaken the study of XY_{2}Z_{2} (C_{2v}) molecules^{g} in the aim of studying the quasi spherical-top SO_{2}F_{2}. Our model has already been used to perform the analysis of the ground state of this molecule. These results have been compared to those of the usual asymmetric-top approach^{h}. We now intend to use this formalism to treat the strongly interacting pentad of this species.

aJ.-P. Champion, M. Loëte and G. Pierre, in Spectroscopy of the Earth’s atmosphere and interstellar medium (K. N. Rao and A.

Weber, Eds.), p. 339–422, Academic Press, San Diego, 1992.

bM. Rotger, V. Boudon and M. Loëte, J. Mol. Spectrosc. , 200, 123–130 (2000).

cM. Rotger, V. Boudon and M. Loëte, J. Mol. Spectrosc. , 200, 131–137 (2000).

dM. Rotger, A. Decrette, V. Boudon M. Loëte, S. Sander and H. Willner, J. Mol. Spectrosc. , 208, 169–179 (2001).

eCh. Wenger, M. Rotger and V. Boudon, J. Quant. Spectrosc. Radiat. Transfer , in press, (2002).

fhttp://www.u-bourgogne.fr/LPUB/c4vTDS.html

gM. Rotger, V. Boudon and M. Loëte, , in preparation, (2002).

hK. Sarka, J. Demaison, L. Margulès, I. Merke, N. Heineking, H. Bürger and H. Ruland, J. Mol. Spectrosc. , 200, 55–64 (2000).